OpenStudy (lgbasallote):

how to turn \[\frac{2s^2 + 1}{s(s+1)^2}\]into partial fractions? special surprise inside

OpenStudy (unklerhaukus):

the normal way

OpenStudy (anonymous):

\[\text{given fraction}=\frac{A}{s}+\frac{B}{s+1}+\frac{C}{(s+1)^2}\]

OpenStudy (lgbasallote):

i got something like A = 1 and C = -3 but i got stuck with B

OpenStudy (lgbasallote):

is root sub applicable? or i can only use gaussian method?

OpenStudy (anonymous):

see attachments.

OpenStudy (lgbasallote):

ahh so you did use gaussian

OpenStudy (anonymous):

I not aware of those names, but yeah..

OpenStudy (anonymous):

*am not.

OpenStudy (lgbasallote):

follow up question... \[\large \mathcal L^{-1}\left \{ \frac{2s^2 + 1}{(s+1)^2}\right \} = 1 + e^{-t} - 3te^{-t}\] is that right?

OpenStudy (anonymous):

I'm sorry, but this is out of my league. I don't even know what that 'L' thing means.

OpenStudy (lgbasallote):

haha =)) yeahhh...good stuffs >:))

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

I don't see the "surprise".

OpenStudy (lgbasallote):

the "surprise" is it's an easy question with a hard question follow up =))

OpenStudy (anonymous):

aha, i see..